# A YTD correspondence for constant scalar curvature metrics

### v1

@@ 8 Algebraic interpretation of $K^{\beta}$-stability
It is known from [71, 18] (see also [61, Proposition 2.38]) that
$\mathcal{J}(\mathcal{X},\mathcal{L})=-\frac{\bar{S}\overline{\mathcal{L}}^{n+1}}{(n+1)V}-\frac{K_{X}\cdot\overline{\mathcal{L}}^{n}}{V}.$ (42)
Here the intersection numbers are taken after a trivial compactification $(\overline{\mathcal{X}},\overline{\mathcal{L}})$ of $(\mathcal{X},\mathcal{L})$ over $\infty\in{\mathbb{P}}^{1}$ and then pull back all the line bundles in question to a common normal model.
In what follows we will give a purely algebraic formula for $Ent^{\beta}(\mathcal{X},\mathcal{L})$ as well, in terms of log canonical thresholds and intersection numbers. By Theorem 7.1,
@@
where the sup is over all geodesic rays $\{v_{t}\}$ arising from semi-ample test-configurations. Let us therefore take an auxiliary semi-ample test configuration $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$, with $\{v_{t}\}$ being the associated geodesic ray.
We know from [71, 18] that
$\beta(I\{v_{t}\}-I\{u_{t}\})=\frac{\beta((\overline{\mathcal{L}^{\prime}})^{n+1}-\overline{\mathcal{L}}^{n+1})}{(n+1)V}.$ (44)
Therefore, it remains to deal with the term $L^{\beta}(\{v_{t}\},\{u_{t}\})$. We will express it in terms of certain log canonical threshold, as shown in the next result. This recovers Berman’s formula for the Ding invariant in the Fano case [5, (3.19)], by taking $(\mathcal{X}^{\prime},\mathcal{L}^{\prime}):=(\mathcal{X},(1-\frac{1}{\beta})\mathcal{L}+\frac{1}{\beta}L)$.
@@
$\mathcal{Z}$$\mathcal{X}^{\prime}$$X\times{\mathbb{C}}$$\mathcal{X}$$\scriptstyle{\mu}$$\scriptstyle{\sigma}$$\scriptstyle{\tau}$
There exist $\mathbb{Q}$-Cartier divisors $F$ and $G$ supported on $\mathcal{Z}_{0}$ such that (cf. [79, Propostion 3.10])
@@
(See [17, §1.5] for the definition of log canonical threshold (lct) for a pair.)
Moreover, when $\mathcal{Z}$ is smooth with simple normal crossing (snc) central fiber, we can write
$\mathcal{Z}_{0}=\sum_{i}a_{i}E_{i},\ \ F=\sum_{i}b_{i}E_{i},\ \ G=\sum_{i}c_{i}E_{i},\ \ K_{\mathcal{Z}/X\times{\mathbb{C}}}:=K_{\mathcal{Z}}-\tau^{*}K_{X\times{\mathbb{C}}}=\sum_{i}d_{i}E_{i}.$
@@
$L^{\beta}(\{v_{t}\},\{u_{t}\})=\sup\left\{\tau\in\mathbb{R}:\int_{X\times\Delta^{*}}e^{\beta(V_{z}-U_{z})-(1+\tau)\log|z|^{2}}\omega^{n}\wedge idz\wedge d\bar{z}<\infty\right\}.$
@@
Indeed, when $\mathcal{Z}$ is smooth with snc central fiber, the complexified geodesic rays $z\mapsto U_{z},\ V_{z}$ when pulled back to $\mathcal{Z}$ have only divisorial singularities along $F$ and $G$, respectively (see [79, Lemma 4.6] and [30, §3]). Then the integrability of $e^{\beta(V_{z}-U_{z})-(1+\tau)\log|z|^{2}}$ is equivalent to requiring that the coefficients of the divisor
$\beta(F-G)-K_{\mathcal{Z}/X\times\mathbb{C}}+(1+\tau)\mathcal{Z}_{0}=\sum_{i}\big(\beta(b_{i}-c_{i})+a_{i}(1+\tau)-d_{i}\big)E_{i}$
are all strictly less than $1$, giving (47).
Lastly, by the pull back formula of pairs (see [17, §1.5]), for a general normal model $\mathcal{Z}$ dominating $\mathcal{X}$, $X\times{\mathbb{C}}$ and $\mathcal{X}^{\prime}$, the quantity $L^{\beta}(\{v_{t}\},\{u_{t}\})+1$ can then be expressed as the lct of the pair $(\mathcal{Z},\ \beta(F-G)-K_{\mathcal{Z}/X\times\mathbb{C}})$ with respect to $\mathcal{Z}_{0}$, giving (46). So we conclude. ∎
Let $\mathcal{T}=(\mathcal{X},\mathcal{L})$ be a semi-ample test configuration. Let $Div_{\mathcal{X}}$ be the set of $\mathbb{Q}$-Cartier divisors $G$ contained in the central fiber $\mathcal{Z}_{0}$ of a $\mathbb{C}^{*}$-equivariant smooth model $\pi:\mathcal{Z}\to\mathcal{X}$, with $\pi$ bijective on $\mathcal{X}\setminus\mathcal{X}_{0}$, and $\mathcal{L}_{G}:=\pi^{*}p_{1}^{*}L-G$ relatively semi-ample on $\mathcal{Z}$.
Thus defined, $(\mathcal{Z},\mathcal{L}_{G})$ is also a semi-ample test configurations, and we define the corresponding divisor $\mathcal{D}_{G}$ as in (45):
$\mathcal{D}_{G}:=\beta(F-G)-K_{\mathcal{Z}}+\tau^{*}K_{X\times{\mathbb{C}}}.$ (48)
@@
Let $\mathcal{T}:=(\mathcal{X},\mathcal{L})$ be a semi-ample test configuration. With the notation introduced above
$K^{\beta}_{\mathcal{T}}=\sup_{G\in Div_{\mathcal{X}}}\Big(\mathrm{lct}_{(\mathcal{\mathcal{Z}},\mathcal{D}_{G})}(\mathcal{Z}_{0})-1+\frac{\beta\overline{\mathcal{L}_{G}}^{n+1}-\beta\overline{\mathcal{L}}^{n+1}}{(n+1)V}\Big)+\frac{\bar{S}\overline{\mathcal{L}}^{n+1}}{(n+1)V}+\frac{K_{X}\cdot\overline{\mathcal{L}}^{n}}{V}.$ (49)
@@
It might be a bit too optimistic to expect that the supremum in (49) is achieved by some $G\in Div_{\mathcal{X}}$. But using a different argument, an intersection type formula for $K^{\beta}_{\mathcal{T}}$ can still be obtained, as will be proved in a sequel paper [43]:
$K^{\beta}_{\mathcal{T}}=\frac{\beta\langle\overline{\mathcal{L}_{\beta}}^{n+1}\rangle-\beta\overline{\mathcal{L}}^{n+1}}{(n+1)V}+\frac{\bar{S}\overline{\mathcal{L}}^{n+1}}{(n+1)V}+\frac{K_{X}\cdot\overline{\mathcal{L}}^{n}}{V},$ (50)
where $\langle\overline{\mathcal{L}_{\beta}}^{n+1}\rangle$ is the top movable intersection product of the (relatively) big line bundle $\overline{\mathcal{L}_{\beta}}:=\overline{\mathcal{L}}+\frac{1}{\beta}\big(K_{\overline{\mathcal{X}}}-\pi^{*}K_{X\times\mathbb{P}^{1}}-\mathcal{X}_{0,red}+\mathcal{X}_{0}\big)$.
@@
When $(X,L)=(X,-K_{X})$ is Fano, our $K^{\beta}(\mathcal{X},\mathcal{L})=K^{\beta}_{\mathcal{T}}$ is directly related to $\mathrm{Ding}(\mathcal{X},\mathcal{L})$, the Ding invariant introduced in [5, 10]. Indeed, choosing $G:=(1-\frac{1}{\beta})F$ in (48) (i.e., taking $\mathcal{L}_{G}:=(1-\frac{1}{\beta})\mathcal{L}+\frac{1}{\beta}(-K_{X})$), one then has (see [90, (2.18)]) $\mathrm{Ding}(\mathcal{X},\mathcal{L})=\mathrm{lct}_{(\mathcal{\mathcal{Z}},\mathcal{D}_{G})}(\mathcal{Z}_{0})-1-\frac{\overline{\mathcal{L}}^{n+1}}{(n+1)V}.$ Thus, (49) implies that (using $\bar{S}=n$ here)
$\begin{split}K^{\beta}(\mathcal{X},\mathcal{L})&\geq\mathrm{Ding}(\mathcal{X},\mathcal{L})+\frac{\beta(\overline{\mathcal{L}_{G}}^{n+1}-\overline{\mathcal{L}}^{n+1})}{(n+1)V}+\frac{(\overline{\mathcal{L}}-(-K_{X}))\cdot\overline{\mathcal{L}}^{n}}{V}\\ &\geq\mathrm{Ding}(\mathcal{X},\mathcal{L})+\frac{\beta}{V}(\overline{\mathcal{L}_{G}}-\overline{\mathcal{L}})\cdot\overline{\mathcal{L}_{G}}^{n}+\frac{(\overline{\mathcal{L}}-(-K_{X}))\cdot\overline{\mathcal{L}}^{n}}{V}\\ &=\mathrm{Ding}(\mathcal{X},\mathcal{L})+\frac{1}{V}(-K_{X}-\overline{\mathcal{L}})\cdot(\overline{\mathcal{L}_{G}}^{n}-\overline{\mathcal{L}}^{n}),\end{split}$
where in the second line we used the Hodge index type estimate [90, Lemma 1.74] (cf. also (2.4) therein) and the last line follows from the relation $\beta(\overline{\mathcal{L}_{G}}-\overline{\mathcal{L}})=-K_{X}$-$\overline{\mathcal{L}}$.
We further argue that the error term $\frac{1}{V}(-K_{X}-\overline{\mathcal{L}})\cdot(\overline{\mathcal{L}_{G}}^{n}-\overline{\mathcal{L}}^{n})$ appearing above conveniently controlled from below when $\beta\gg 1$. Indeed, up to adding some multiple $\mathcal{X}_{0}$ to $\mathcal{L}$, we can assume that $(-K_{X}-\overline{\mathcal{L}})$, when pulled back to a common normal model $\mathcal{Z}$, is an effective $\mathbb{Q}$-divisor supported in $\mathcal{Z}_{0}$ and it does not contain any multiple of $\mathcal{Z}_{0}$ (so that $\overline{\mathcal{L}}\cdot(-K_{X})^{n}=0$). Then using that $\overline{\mathcal{L}_{G}}-(1-\frac{1}{\beta})\overline{\mathcal{L}}$ is nef, we deduce that
$\begin{split}\frac{1}{V}(-K_{X}-\overline{\mathcal{L}})\cdot(\overline{\mathcal{L}_{G}}^{n}-\overline{\mathcal{L}}^{n})\geq-\frac{(1-(1-\beta^{-1})^{n})}{V}(-K_{X}-\overline{\mathcal{L}})\cdot\overline{\mathcal{L}}^{n}=-\varepsilon_{\beta}\mathbf{I}(\mathcal{X},\mathcal{L}),\end{split}$
where $\varepsilon_{\beta}:=(1-(1-\beta^{-1})^{n})$ and $\mathbf{I}(\mathcal{X},\mathcal{L})$ denotes the $\mathbf{I}$-norm of $(\mathcal{X},\mathcal{L})$ (see [90, (2.2)]). Using the relation $\mathbf{I}\leq(n+1)\mathbf{J}$ [90, Proposition 2.9], we finally arrive at
@@
9 Connections with the non-Archimedean point of view
We briefly point out the relationship of our work with the non-Archimedean point of view and give an alternative argument of a recent remarkable result due to Boucksom–Jonsson [23]. For background and terminology on non-Archimedean geometry we refer to [17].
Assume that $(X,L)$ is a polarized manifold with $\omega\in c_{1}(L)$. Let us first give a non-Archimedean interpretation for the radial $\beta$-entropy. In light of Theorem 7.1, it suffices to give a non-Archimedean formula for $L^{\beta}(\{v_{t}\},\{u_{t}\})$. We achieve this by adapting the treatment of the Ding energy from [10, Theorem 5.4].
@@ 9 Connections with the non-Archimedean point of view
Let $\{u_{t}\}$ be a bounded geodesic ray, and let $\{v_{t}\}$ be a geodesic ray arising from a semi-ample test configuration. Then for any $\beta>0$,
@@
Here $u^{an}$ and $v^{an}$ are non-Archimedean potentials associated with $\{u_{t}\}$ and $\{v_{t}\}$ in the sense of [10, Definition 4.2] (see also [41, Proposition 3.1]). Moreover, $X^{div}_{\mathbb{Q}}$ denotes the $\mathbb{Q}$-valued divisorial valuations on $X$ (including the trivial one).
@@
Up to further subtracting $Ct$ from $\{v_{t}\}$ we can assume that $e^{\beta(V_{z}-U_{z})}\in L^{1}(X\times\Delta)$. Put for simplicity $\mu:=e^{\beta(V_{z}-U_{z})}\omega^{n}\wedge idz\wedge d\bar{z}$. Then, in terms of the complex singularity exponent (see [42, §2.2]), we can write
@@
Noting that $V_{z}$ has algebraic singularities (recall (14)) and that $V_{z}$, $U_{z}$ and $\log|z|^{2}$ are all $S^{1}$-invariant and bounded away from $X\times\{0\}$, we can apply the valuative criterion [10, Theorem B.7] in the same way as in the short proof of [42, Theorem 2.3] to deduce that
@@
$A_{X\times\Delta}(E)+\beta\nu(V,E)-\beta\nu(U,E)\geq(1+\tau)\nu(\log|z|^{2},E)$ (51)
@@
$b_{E}:=\nu(\log|z|^{2},E).$
By [17, Theorem 4.6], all divisorial valuations $w\in X^{div}_{\mathbb{Q}}$ (including the trivial valuation) are in one-to-one correspondence with $\frac{1}{b_{E}}\mathrm{ord}_{E}$, via the Gauss map/extension. In fact, the Gauss extension $G(w)$ of any $w\in X^{div}_{\mathbb{Q}}$ is of the form $G(w)=\frac{1}{b_{E}}\mathrm{ord}_{E}$ for some $E$ over $X\times\{0\}$ and we have the relation $A_{X}(w)=\frac{1}{b_{E}}A_{X\times\Delta}(E)-1$ (by [21, Theorem A.10]). Moreover, we have that (cf. [10, Definition 4.2])
$u^{an}(w)=-\frac{1}{b_{E}}\nu(U,E)\text{ and }v^{an}(w)=-\frac{1}{b_{E}}\nu(V,E).$
Revisiting (51), we obtain that $L^{\beta}(\{v_{t}\},\{u_{t}\})$ is equal to the largest $\tau\in\mathbb{R}$ such that
@@
Let $\{u_{t}\},\{v_{t}\}$ be geodesic rays induced by a model filtration and a semi ample test configuration respectively. Then for any $\beta>0$:
@@
where $L^{NA}(f):=\inf_{w\in X^{div}}(A_{X}(w)+f(w))$ is the natural non-Archimedean functional from [20, §2.6], defined for continuous non-Archimedean functions.
@@
Since model filtrations are linearly bounded, the associated ray $\{u_{t}\}$ must be bounded [76], hence the previous result is applicable. Both $v^{an}$ and $u^{an}$ are continuous on $X^{an}$ by the sentence below [61, formula (28)], hence (52) holds, by the definition of $L^{NA}$. ∎
@@
Here $\mathrm{MA}(u^{an})$ denotes the non-Archimedean Monge-Ampère measure of $u^{an}$ and $Ent^{NA}$ denotes the non-Archimedean entropy (see [17, §6.7 and Definition 7.17]).
@@
Due to Theorem 5.2, [61, Proposition 6.3] and [75, Theorem 4.4.1], we can approximate both sides of (53) using a sequence of geodesic rays arising from model filtrations in the sense of [61, Definition 2.7].
So in what follows we simply assume that $\{u_{t}\}$ itself a geodesic ray arising from a model filtration. Pick any geodesic ray $\{v_{t}\}$ arising from a semi-ample test configuration. Corollary 9.2 yields $L^{\beta}(\{v_{t}\},\{u_{t}\})=L^{NA}(\beta({u}^{an}-v^{an})).$ Applying [20, Proposition 2.8] and [22, (4.5)],

### v2

@@ 7 Threshold type interpretation of $K^{\beta}$-stability
It is known from [81, 22] (see also [70, Proposition 2.38]) that
$\mathcal{J}(\mathcal{X},\mathcal{L})=-\frac{\bar{S}\mathcal{L}^{n+1}}{(n+1)V}-\frac{K_{X}\cdot\mathcal{L}^{n}}{V}.$ (56)
Here the intersection numbers are taken after a trivial compactification of $(\mathcal{X},\mathcal{L})$ over $\infty\in{\mathbb{P}}^{1}$ and then pull back all the line bundles in question to a common normal model. In what follows we will provide a purely algebraic description for $Ent^{\beta}(\mathcal{X},\mathcal{L})$ as well, in terms of log canonical thresholds and intersection numbers.
By Theorem 6.1, for any $\beta>1$,
@@
where the sup is over all geodesic rays $\{v_{t}\}$ arising from test-configurations. Let us therefore take an auxiliary test configuration $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$, with $\{v_{t}\}$ being the associated geodesic ray.
We know from the slope formulas in [81, 22] that
$\beta(I\{v_{t}\}-I\{u_{t}\})=\frac{\beta(({\mathcal{L}^{\prime}})^{n+1}-\mathcal{L}^{n+1})}{(n+1)V}.$ (58)
Therefore, it remains to deal with the term $L^{\beta}(\{v_{t}\},\{u_{t}\})$. We will express it in terms of certain log canonical threshold, as shown in the next result. It recovers Berman’s formula for the Ding invariant in the Fano case [4, (3.19)], by taking $(\mathcal{X}^{\prime},\mathcal{L}^{\prime}):=(\mathcal{X},(1-\frac{1}{\beta})\mathcal{L}+\frac{1}{\beta}L)$.
@@
$\mathcal{Z}$$\mathcal{X}^{\prime}$$X\times{\mathbb{P}}^{1}$$\mathcal{X}$$\scriptstyle{\mu}$$\scriptstyle{\sigma}$$\scriptstyle{\tau}$
There exist $\mathbb{Q}$-Cartier divisors $F$ and $G$ supported on $\mathcal{Z}_{0}$ such that (cf. [90, Propostion 3.10])
@@
(See [21, §1.5] for the definition of log canonical threshold (lct) for a pair.)
Moreover, when $\mathcal{Z}$ is smooth with simple normal crossing (snc) central fiber, write
$\mathcal{Z}_{0}=\sum_{i}a_{i}E_{i},\ \ F=\sum_{i}b_{i}E_{i},\ \ G=\sum_{i}c_{i}E_{i},\ \ K_{\mathcal{Z}}-\tau^{*}K_{X\times{\mathbb{P}}^{1}}=\sum_{i}d_{i}E_{i}.$
@@
$L^{\beta}(\{v_{t}\},\{u_{t}\})=\sup\left\{\tau\in\mathbb{R}:\int_{X\times\Delta^{*}}e^{\beta(V_{z}-U_{z})-(1+\tau)\log|z|^{2}}\omega^{n}\wedge idz\wedge d\bar{z}<\infty\right\}.$ (62)
@@
Indeed, when $\mathcal{Z}$ is smooth with snc central fiber, the complexified geodesic rays $z\mapsto U_{z},\ V_{z}$ when pulled back to $\mathcal{Z}$ is $L^{\infty}$ compatible with the divisors $F$ and $G$, respectively (recall (21)). Then the integrability of $e^{\beta(V_{z}-U_{z})-(1+\tau)\log|z|^{2}}$ is equivalent to requiring that the coefficients of the divisor $\beta(F-G)-K_{\mathcal{Z}/X\times{\mathbb{P}}^{1}}+(1+\tau)\mathcal{Z}_{0}=\sum_{i}\big(\beta(b_{i}-c_{i})+a_{i}(1+\tau)-d_{i}\big)E_{i}$ are all strictly less than 11, giving (61).
Lastly, by the pull back formula of pairs (see [21, §1.5]), for a general normal model $\mathcal{Z}$ dominating $\mathcal{X}$, $X\times{\mathbb{P}}^{1}$ and $\mathcal{X}^{\prime}$, the quantity $L^{\beta}(\{v_{t}\},\{u_{t}\})+1$ can then be expressed as the lct of the pair $(\mathcal{Z},\ \beta(F-G)-K_{\mathcal{Z}/X\times\mathbb{C}})$ with respect to $\mathcal{Z}_{0}$, giving (60). So we conclude. ∎
Let $\mathcal{T}=(\mathcal{X},\mathcal{L})$ be a test configuration dominating $X\times{\mathbb{P}}^{1}$. Let $Div_{\mathcal{X}}$ be the set of $\mathbb{Q}$-Cartier divisors $G$ contained in the central fiber $\mathcal{Z}_{0}$ of a $\mathbb{C}^{*}$-equivariant smooth model $\pi:\mathcal{Z}\to\mathcal{X}$, with $\pi$ bijective on $\mathcal{X}\setminus\mathcal{X}_{0}$, and $\mathcal{L}_{G}:=\pi^{*}p_{1}^{*}L-G$ relatively semi-ample on $\mathcal{Z}$.
Thus defined, $(\mathcal{Z},\mathcal{L}_{G})$ is also a test configuration of $(X,L)$, and we define the corresponding divisor $\mathcal{D}_{G}$ as in (59):
$\mathcal{D}_{G}:=\beta(F-G)-K_{\mathcal{Z}}+\tau^{*}K_{X\times{\mathbb{P}}^{1}}.$ (63)
@@
Let $\mathcal{T}:=(\mathcal{X},\mathcal{L})$ be a test configuration dominating $X\times{\mathbb{P}}^{1}$. With the notation introduced above, for any $\beta>1$ we have
$K^{\beta}_{\mathcal{T}}=\sup_{G\in Div_{\mathcal{X}}}\Big(\mathrm{lct}_{(\mathcal{\mathcal{Z}},\mathcal{D}_{G})}(\mathcal{Z}_{0})-1+\frac{\beta({\mathcal{L}_{G}}^{n+1}-{\mathcal{L}}^{n+1})}{(n+1)V}\Big)+\frac{\bar{S}\mathcal{L}^{n+1}}{(n+1)V}+\frac{K_{X}\cdot\mathcal{L}^{n}}{V}.$ (64)
@@
When $(X,L)=(X,-K_{X})$ is Fano, our $K^{\beta}(\mathcal{X},\mathcal{L})$ is directly related to $\mathrm{Ding}(\mathcal{X},\mathcal{L})$, the Ding invariant introduced in [4, 9]. Indeed, choosing $G:=(1-\frac{1}{\beta})F$ in (63) (i.e., taking $\mathcal{L}_{G}:=(1-\frac{1}{\beta})\mathcal{L}+\frac{1}{\beta}(-K_{X})$), one then has (by [101, (2.18)]) $\mathrm{Ding}(\mathcal{X},\mathcal{L})=\mathrm{lct}_{(\mathcal{\mathcal{Z}},\mathcal{D}_{G})}(\mathcal{Z}_{0})-1-\frac{\mathcal{L}^{n+1}}{(n+1)V}.$ Thus, (64) implies that (using $\bar{S}=n$ here)
$\begin{split}K^{\beta}(\mathcal{X},\mathcal{L})&\geq\mathrm{Ding}(\mathcal{X},\mathcal{L})+\frac{\beta({\mathcal{L}_{G}}^{n+1}-{\mathcal{L}}^{n+1})}{(n+1)V}+\frac{({\mathcal{L}}-(-K_{X}))\cdot{\mathcal{L}}^{n}}{V}\\ &\geq\mathrm{Ding}(\mathcal{X},\mathcal{L})+\frac{\beta}{V}({\mathcal{L}_{G}}-{\mathcal{L}})\cdot{\mathcal{L}_{G}}^{n}+\frac{({\mathcal{L}}-(-K_{X}))\cdot{\mathcal{L}}^{n}}{V}\\ &=\mathrm{Ding}(\mathcal{X},\mathcal{L})+\frac{1}{V}(-K_{X}-{\mathcal{L}})\cdot({\mathcal{L}_{G}}^{n}-{\mathcal{L}}^{n}),\end{split}$
where in the second line we used the Hodge index type estimate [101, Lemma 1.74] (cf. also (2.4) therein) and the last line follows from the relation $\beta({\mathcal{L}_{G}}-\overline{\mathcal{L}})=-K_{X}$-$\mathcal{L}$}.
We further argue that the error term $\frac{1}{V}(-K_{X}-{\mathcal{L}})\cdot({\mathcal{L}_{G}}^{n}-{\mathcal{L}}^{n})$ appearing above can be conveniently controlled from below when $\beta\gg 1$. Indeed, up to adding some multiple $\mathcal{X}_{0}$ to $\mathcal{L}$, we can assume that $(-K_{X}-{\mathcal{L}})$, when pulled back to a common normal model $\mathcal{Z}$, is an effective $\mathbb{Q}$-divisor supported in $\mathcal{Z}_{0}$ and it does not contain any multiple of $\mathcal{Z}_{0}$ (so that $\overline{\mathcal{L}}\cdot(-K_{X})^{n}=0$). Then using that ${\mathcal{L}_{G}}-(1-\frac{1}{\beta}){\mathcal{L}}$ is nef, we deduce that
$\begin{split}\frac{1}{V}(-K_{X}-{\mathcal{L}})\cdot({\mathcal{L}_{G}}^{n}-{\mathcal{L}}^{n})\geq-\frac{(1-(1-\beta^{-1})^{n})}{V}(-K_{X}-{\mathcal{L}})\cdot{\mathcal{L}}^{n}=-\varepsilon_{\beta}\mathbf{I}(\mathcal{X},\mathcal{L}),\end{split}$
where $\varepsilon_{\beta}:=(1-(1-\beta^{-1})^{n})$ and $\mathbf{I}(\mathcal{X},\mathcal{L})$ denotes the $\mathbf{I}$-norm of $(\mathcal{X},\mathcal{L})$ (see [101, (2.2)]). Using the relation $\mathbf{I}\leq(n+1)\mathbf{J}$ [101, Proposition 2.9], we finally arrive at
@@
Beyond the Fano setting, we hope that (64) will lead to new valuative criterions for the existence of cscK metrics, generalizing those for KE metrics (cf. [92, 68, 58, 59, 15, 1, 2, 101]).
8 Connections with non-Archimedean geometry
We point out the relationship of our work with the non-Archimedean point of view and give an alternative argument of a recent remarkable result of Boucksom–Jonsson [27].
Assume that $(X,L)$ is a polarized manifold with $\omega\in c_{1}(L)$. Let us first give a non-Archimedean interpretation for the radial $\beta$-entropy. In light of Theorem 6.1, it suffices to give a non-Archimedean formula for $L^{\beta}(\{v_{t}\},\{u_{t}\})$. We achieve this by adapting the treatment for the Ding energy in [9, Theorem 5.4].
@@ 8 Connections with non-Archimedean geometry
Let $\{u_{t}\}$ be a bounded geodesic ray, and let $\{v_{t}\}$ be a geodesic ray arising from a test configuration. Then for any $\beta>0$,
@@
Here $u^{an}$ and $v^{an}$ are non-Archimedean potentials associated with $\{u_{t}\}$ and $\{v_{t}\}$ in the sense of [9, Definition 4.2] (see also [45, Proposition 3.1]). Moreover, $X^{div}_{\mathbb{Q}}$ denotes the $\mathbb{Q}$-valued divisorial valuations on $X$ (including the trivial one).
@@
Up to further subtracting $Ct$ from $\{v_{t}\}$ we can assume that $e^{\beta(V_{z}-U_{z})}\in L^{1}(X\times\Delta)$. Put for simplicity $\mu:=e^{\beta(V_{z}-U_{z})}\omega^{n}\wedge idz\wedge d\bar{z}$. Then, in terms of the complex singularity exponent (see [46, §2.2]), we can write
@@
Noting that $V_{z}$ has algebraic singularities (recall (19)) and that $V_{z}$, $U_{z}$ and $\log|z|^{2}$ are all $S^{1}$-invariant and bounded away from $X\times\{0\}$, we can apply the valuative criterion [9, Theorem B.7] in the same way as in the short proof of [46, Theorem 2.3] to deduce that
@@
$A_{X\times\Delta}(E)+\beta\nu(V_{z},E)-\beta\nu(U_{z},E)\geq(1+\tau)\nu(\log|z|^{2},E)$ (65)
@@
Let $\{u_{t}\},\{v_{t}\}$ be geodesic rays induced by a model and a test configuration, respectively. Then for any $\beta>0$:
@@
where $L^{NA}(f):=\inf_{w\in X^{div}}(A_{X}(w)+f(w))$ is the natural non-Archimedean functional from [23, §2.6], defined for continuous non-Archimedean functions.
@@
Since $\{u_{t}\}$ is induced by the filtration in [70, Definition 2.7], it must be bounded [86, §9], hence the previous result is applicable. Both $v^{an}$ and $u^{an}$ are continuous on $X^{an}$ by the sentence below [70, formula (28)], hence (66) holds by the definition of $L^{NA}$. ∎
@@
Here $\mathrm{MA}(u^{an})$ denotes the non-Archimedean Monge–Ampère measure of $u^{an}$ and $Ent^{NA}$ denotes the non-Archimedean entropy (see [21, §6.7 and Definition 7.17]).
@@
Due to Theorem 5.2, [70, Proposition 6.3] and [85, Theorem 4.4.1], we can approximate both sides of (67) using a sequence of geodesic rays arising from models. So in what follows we simply assume that $\{u_{t}\}$ itself a geodesic ray arising from a model . Pick any geodesic ray $\{v_{t}\}$ arising from a test configuration. Corollary 8.2 yields $L^{\beta}(\{v_{t}\},\{u_{t}\})=L^{NA}(\beta({u}^{an}-v^{an})).$ Applying [23, Proposition 2.8] and [26, (4.5)],